Consider the function f(x, y) = x^3 − 2xy + y^2 + 5.

(a) Find the equation for the tangent plane to the graph of z = f(x, y) at the point (2, 3, f(2, 3)).

(b) Calculate an estimate for the value f(2.1, 2.9) using the standard linear approximation of f at (2, 3).

(c) Find the normal line to the zero level surface of F(x, y, z) = f(x, y) − z at the point (2, 3, f(2, 3))
Answer #1

fracy) = x3 - 2xy + y2 + 5 f (2,3) = 8 - 12 + 9+5 - 10. Rhe s. fxz zurray fx (2, 3) = 6. fy - - 2x +2y fy (2, 3) = 2. Lray) =